Compute Financial ratios:

Price-earnings, Cash Dicidend payot, Debt Ratio, Debt-to-Equity, and Times Interest earned

Orange Company
Income Statement
For the Years Ended December 31
2013 2012
Net sales (all on account) $            600,000 $                520,000
Expenses:
Cost of Goods Sold $            415,000 $                354,000
Selling and administrative $            120,800 $                114,600
Interest Expense $                7,800 $                    6,000
Income Tax Expense $              18,000 $                  14,000
Total expenses $            561,600 $                488,600
Net Income $              38,400 $                  31,400
Additional Data:
1. The common stock recently sold at $19.50 per share.
2. Cash dividends in the amount of $15,400 were paid-out in 2013.
Orange Company
Balance Sheets
December 31
2013 2012
Assets
Current Assets
Cash $        21,000 $        18,000
Short-term investments $        18,000 $        15,000
Accounts Receivable $        86,000 $        74,000
Inventory $        90,000 $        70,000
Total Current Assets $      215,000 $      177,000
Plant Assets $      423,000 $      383,000
Total Assets $      638,000 $      560,000
Liabilities and Stockholder's Equity
Current Liabilities
Accounts Payable $      122,000 $      110,000
Income Taxes Payable $        23,000 $        20,000
Total Curent Liabilities $      145,000 $      130,000
Long-term Liabilities
Bonds Payable $      120,000 $        80,000
Total Liabilities $      265,000 $      210,000
Stockholder's Equity
Common Stock ($5 par value) $      150,000 $      150,000
Retained Earnings $      223,000 $      200,000
Total Stockholder's Equity $      373,000 $      350,000
Total Liabilities and Stockholder's Equity $      638,000 $      560,000
Answer #1

Solution: Calculation of various financial ratios for the year 2013:

1. Price-Earnings ratio = Market value per share / Earning per share = 19.5/1.28 = 15.23

Note: Earning per share = Net Income / common shares = 38400/30000 = $1.28

2. Cash-Dividend Payout ratio = Cash Dividend / Net Income = 15400/38400 = 0.401 or 40.10%

3. Debt ratio = Total liabilities / Total Assets = 265000/638000 = 0.4153 or 41.53%

4. Debt-to-equity ratio = Total liabilities / Total equity = 265000/373000 = 0.7104 or 71.04%

5. Times Interest Earned ratio = EBIT / Interest Expense = (600000-415000-120800) / 7800

= 8.23 times